### v1

# 3DEnhancer: Consistent Multi-View Diffusion for 3D Enhancement

Abstract
Despite advances in neural rendering, due to the scarcity of high-quality 3D datasets and the inherent limitations of multi-view diffusion models, view synthesis and 3D model generation are restricted to low resolutions with suboptimal multi-view consistency. In this study, we present a novel 3D enhancement pipeline, dubbed 3DEnhancer, which employs a multi-view latent diffusion model to enhance coarse 3D inputs while preserving multi-view consistency. Our method includes a pose-aware encoder and a diffusion-based denoiser to refine low-quality multi-view images, along with data augmentation and a multi-view attention module with epipolar aggregation to maintain consistent, high-quality 3D outputs across views. Unlike existing video-based approaches, our model supports seamless multi-view enhancement with improved coherence across diverse viewing angles. Extensive evaluations show that 3DEnhancer significantly outperforms existing methods, boosting both multi-view enhancement and per-instance 3D optimization tasks.
Figure 1: Our proposed 3DEnhancer showcases excellent capabilities in enhancing multi-view images generated by various models. As shown in (a), it can significantly improve texture details, correct texture errors, and enhance consistency across views. Beyond enhancement, as illustrated in (b), 3DEnhancer also supports texture-level editing, including regional inpainting, and adjusting texture enhancement strength via noise level control. (Zoom-in for best view)
1 Introduction
The advancements in generative models [24, 19] and differentiable rendering [43] have paved the way for a new research field known as neural rendering [60]. In addition to pushing the boundaries of view synthesis [32], the generation and editing of 3D models [81, 54, 30, 35, 84, 39, 36, 34, 13] has become achievable. These methods are trained on the large-scale 3D datasets, e.g., Objaverse dataset [14], enabling fast and diverse 3D synthesis.
Despite these advances, several challenges remain in 3D generation. A key limitation is the scarcity of high-quality 3D datasets; unlike the billions of high-resolution image and video datasets available [50], current 3D datasets [15] are limited to a much smaller scale [47]. Another limitation is the dependency on multi-view (MV) diffusion models [54, 53]. Most state-of-the-art 3D generative models [58, 71] follow a two-stage pipeline: first, generating multi-view images conditioned on images or text [65, 54], and then reconstructing 3D models from these generated views [27, 58]. Consequently, the low-quality results and view inconsistency issues of multi-view diffusion models [54] restrict the quality of the final 3D output. Besides, existing novel view synthesis methods [43, 32] usually require dense, high-resolution input views for optimization, making 3D content creation challenging when only low-resolution sparse captures are available.
In this study, we address these challenges by introducing a versatile 3D enhancement framework, dubbed 3DEnhancer, which leverages a text-to-image diffusion model as the 2D generative prior to enhance general coarse 3D inputs. The core of our proposed method is a multi-view latent diffusion model (LDM) [48] designed to enhance coarse 3D inputs while ensuring multi-view consistency. Specifically, the framework consists of a pose-aware image encoder that encodes low-quality multi-view renderings into latent space and a multi-view-based diffusion denoiser that refines the latent features with view-consistent blocks. The enhanced views are then either used as input for multi-view reconstruction or directly serve as reconstruction targets for optimizing the coarse 3D inputs.
To achieve practical results, we introduce diverse degradation augmentations [69] to the input multi-view images, simulating the distribution of coarse 3D data. In addition, we incorporate efficient multi-view row attention [36, 28] to ensure consistency across multi-view features. To further reinforce coherent 3D textures and structures under significant viewpoint changes, we also introduce near-view epipolor aggregation modules, which directly propagate corresponding tokens across near views using epipolar-constrained feature matching [18, 11]. These carefully designed strategies effectively contribute to achieving high-quality, consistent multi-view enhancement.
The most relevant works to our study are 3D enhancement approaches using video diffusion models [52, 74]. While video super-resolution (VSR) models [76] can also be adapted for 3D enhancement, several challenges that make them less suitable for use as generic 3D enhancers. First, these methods are limited to enhancing 3D model reconstructions through per-instance optimization, whereas our approach can seamlessly enhance 3D outputs by integrating multi-view enhancement into the existing two-stage 3D generation frameworks (e.g., from “MVDream [54] $\to$ LGM [58]” to “MVDream $\to$ 3DEnhancer $\to$ LGM”). Second, video models often struggle with long-term consistency and fail to correct generation artifacts in 3D objects under significant viewpoint variations. Besides, video diffusion models based on temporal attention [3] face limitations in handling long videos due to memory and speed constraints. In contrast, our multi-view enhancer models texture correspondences across various views both implicitly and explicitly, by utilizing multi-view raw attention and near-view epipolar aggregation, leading to superior view consistency and higher efficiency.
In summary, we present a novel 3DEnhancer for generic 3D enhancement using multi-view denoising diffusion. Our contributions include a robust data augmentation pipeline, and the hybrid view-consistent blocks that integrate multi-view row attention and a near-view epipolar aggregation module to promote view consistency. Compared to existing enhancement methods, our multi-view 3D enhancement framework is more versatile and supports texture refinement. We conduct extensive experiments on both multi-view enhancement and per-instance optimization tasks to evaluate the model’s components. Our proposed pipeline significantly improves the quality of coarse 3D objects and consistently surpasses existing alternatives.
2 Related Work
3D Generation with Multi-view Diffusion. The success of 2D diffusion models [57, 24] has inspired their application to 3D generation. Score distillation sampling [46, 70] distills 3D from a 2D diffusion model but faces challenges like expensive optimization, mode collapse, and the Janus problem. More recent methods propose learning the 3D via a two-stage pipeline: multi-view images generation [54, 42, 53, 71] and feed-forward 3D reconstruction [27, 75, 58]. Though yielding promising results, their performance is bounded by the quality of the multi-view generative models, including the violation of strict view consistency [39] and failing to scale up to higher resolution [53]. Recent work has focused on developing more 3D-aware attention operations, such as epipolar attention [61, 28] and row-wise attention [36]. However, we find that enforcing strict view consistency remains challenging when relying solely on attention-based operations.
Image and Video Super-Resolution. Image and video SR aims to improve visual quality by upscaling low-resolution content to high resolution. Research in this field has evolved from focusing on pre-defined single degradations [86, 66, 88, 37, 9, 12, 67, 5, 6, 38] (e.g., bicubic downsampling) to addressing unknown and complex degradations [82, 69, 7, 90] in real-world scenarios. To tackle real-world enhancement, some studies [82, 69, 7, 90] introduce effective degradation pipelines that simulate diverse degradations for data augmentation during training, significantly boosting performance in handling real-world cases. To achieve photorealistic enhancement, recent work has integrated various generative priors to produce detailed textures, including StyleGAN [4, 68, 79], codebook [89, 8], and the latest diffusion models [64, 91]. For instance, StableSR [64] leverages the pretrained image diffusion model, i.e., Stable Diffusion (SD) [48], for image enhancement, while Upscale-A-Video [91] further extends the diffusion model for video upscaling. Video SR networks commonly employ recurrent frame fusion [87, 67], optical flow-guided propagation [5, 6, 7, 38] or temporal attention [91] to enhance temporal consistency across adjacent frames. However, due to large spatial misalignments from viewpoint changes, these methods face challenges in establishing long-range correspondences across multi-view images, making them unsuitable for multi-view fusion for 3D. In this study, we focus on exploiting a text-to-image diffusion model to achieve robust 3D enhancement while preserving view consistency.
3D Texture Enhancement. With the rapid advancement of 3D generative models [84, 2, 35, 34, 13], attention is paid to further improve 3D generation quality through a cascade 3D enhancement module. Meta 3D Gen [2, 1] proposes a UV space enhancement model to achieve sharper textures. However, training the UV-specific enhancement model requires spatially continuous UV maps, which are limited in both quantities [14] and qualities [29]. Intex [59] and SyncMVD [40] also employ UV space for generating and enhancing 3D textures. However, these techniques are specifically designed for 3D mesh with UV coordinates, making them unsuitable for other 3D representations like 3DGS [32]. Unique3D [72] and CLAY [84] apply 2D enhancement module RealESRGAN [69] directly to the generated multi-view outputs. Though straightforward, this approach risks compromising 3D consistency across the multi-view results. MagicBoost [78] introduces a 3D refinement pipeline but relies on computationally expensive SDS optimization. SuperGaussian [52] and 3DGS-Enhancer [74] propose to enhance 3D through 2D video generative priors [3, 76]. These pre-trained video models struggle to maintain long-range consistency under large viewpoint variations, making them less effective at fixing texture errors in multi-view generation.
3 Methodology
Figure 2: An overview of 3DEnhancer. By harnessing the power of generative priors, 3DEnhancer adapts a text-to-image diffusion model to a multi-view framework aimed at 3D enhancement. It is compatible with multi-view images generated by models like MVDream [54] or those rendered from coarse 3D representations, such as NeRFs [43] and 3DGS [32]. Given low-quality multi-view images along with their corresponding camera poses, 3DEnhancer aggregates multi-view information within a DiT [44] framework using row attention and epipolar aggregation modules, improving visual quality while preserving high consistency across views. Furthermore, the model supports texture-level editing via text prompts and adjustable noise levels, allowing users to correct texture errors and control the enhancement strength.
A common pipeline in current 3D generation involves an image-to-multiview stage [65], followed by multiview-to-3D [58] generation that converts these multi-view images into a 3D object. However, due to limitations in resolution and view consistency [39], the resulting 3D outputs often lack high-quality textures and detailed geometry. The proposed multi-view enhancement network, 3DEnhancer, aims at improving the quality of 3D representations. Our motivation is that if we can obtain high-quality and view-consistent multi-view images, then the quality of 3D generation can be correspondingly enhanced.
As illustrated in Fig. 2, our framework employs a Diffusion Transformer (DiT) based LDM [44, 10] as the backbone. We incorporate a pose-aware encoder and view-consistent DiT blocks to ensure multi-view consistency, allowing us to leverage the powerful multi-view diffusion models to enhance both coarse multi-view images and 3D models. The enhanced multi-view images can improve the performance of pre-trained feed-forward 3D reconstruction models, e.g., LGM [58], as well as optimize a coarse 3D model through iterative updates.
Preliminary: Multi-view Diffusion Models. LDM [48, 62, 24] is designed to acquire a prior distribution $p_{\bm{\theta}}({\mathbf{z}})$ within the perceptual latent space, whose training data is the latent obtained from the trained VAE encoder $\mathcal{E}_{\bm{\phi}}$. By training to predict a denoised variant of the noisy input ${\mathbf{z}}_{t}=\alpha_{t}{\mathbf{z}}+\sigma_{t}\bm{\epsilon}$ at each diffusion step $t$, $\bm{\epsilon}_{\Theta}$ gradually learns to denoise from a standard Normal prior $\mathcal{N}(\mathbf{0},\mathbf{I})$ by solving a reverse SDE [24].
Similarly, multi-view diffusion generation models [77, 54] consider the joint distribution of multi-view images $\mathcal{X}=\{{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{N}\}$, where each set of $\mathcal{X}$ contains RGB renderings ${\mathbf{x}}_{{\mathbf{v}}}\in\mathbb{R}^{H\times W\times 3}$ from the same 3D asset given viewpoints $\mathcal{C}=\{\pi_{1},\ldots,\pi_{N}\}$. The latent diffusion process is identical to diffusing each encoded latent ${\mathbf{z}}=\mathcal{E}_{\bm{\phi}}({\mathbf{x}})$ independently with the shared noise schedule: $\mathcal{Z}_{t}=\{\alpha_{t}{\mathbf{z}}+\sigma_{t}\bm{\epsilon}\mid{\mathbf{z}}\in\mathcal{Z}\}$. Formally, given the multi-view data $\mathcal{D}_{mv}:=\{\mathcal{X},\mathcal{C},y\}$, the corresponding diffusion loss is defined as:
$\mathcal{L}_{MV}(\theta,\mathcal{D}_{mv})=\mathbb{E}_{\mathcal{Z},y,\pi,t,\bm{\epsilon}}\left[\|\epsilon-\bm{\epsilon}_{\Theta}(\mathcal{Z}_{t};y,\pi,t)\|_{2}^{2}\right],$ (1)
where $y$ is the optional text or image condition.
3.1 Pose-aware Encoder
Given the posed multi-view images $\mathcal{X}$, we add controllable noise to the images as an augmentation to enable controllable refinement, as described later in Sec. 3.3. To further inject camera condition for each view $v$, we follow the prior work [55, 77, 35, 58], and concatenate Plücker coordinates ${\mathbf{r}}_{\mathbf{v}}^{i}=({\mathbf{d}}^{i},{\mathbf{o}}^{i}\times{\mathbf{d}}^{i})\in\mathbb{R}^{6}$ with image RGB values ${\mathbf{x}}_{{\mathbf{v}}}^{i}\in\mathbb{R}^{3}$ along the channel dimension. Here, ${\mathbf{o}}^{i}$ and ${\mathbf{d}}^{i}$ are the ray origin and ray direction for pixel $i$ from view ${\mathbf{v}}$, and $\times$ denotes the cross product. We then send the concatenated results to a trainable pose-aware multi-view encoder $\mathcal{E}_{\bm{\psi}}$, whose outputs are injected into the pre-trained DiT through a learnable copy [83]. The visual illustration of our design is shown in Fig. 2.
3.2 View-Consistent DiT Block
The main challenge of 3D enhancement is achieving precise view consistency across generated 2D multi-view images. Multi-view diffusion methods commonly rely on multi-view attention layers to exchange information across different views, aiming to generate multiview-consistent images. A prevalent approach is extending self-attention to all views, known as dense multi-view attention [54, 42]. While effective, this method significantly raises both computational demands and memory requirements.
To further enhance the effectiveness and efficiency of inter-view aggregation, we introduce two efficient modules in the DiT blocks: multi-view row attention and near-view epipolar aggregation, as shown in Fig. 2.
Multi-view Row Attention. To enhance the noisy input views to higher resolution, e.g., $512\times 512$, efficient multi-view attention is required to facilitate cross-view information fusion. Considering the epipolar constraints [21], the 3D correspondences across multiple views always lie on the epipolar line [61, 28]. Since our diffusion denoising is performed on 16×16$\times$16 × downsampled features [10], and typical multi-view settings often involve elevation angles around $0^{\circ}$, we assume that horizontal rows approximate the epipolar line. Therefore, we adopt the special epipolar attention, specifically the multi-view row attention [36], which enables efficient information exchange among multi-view features.
Specifically, the input cameras are chosen to look at the object with their $Y$ axis aligned with the gravity direction and cameras’ viewing directions are approximately horizontal (i.e., the pitch angle is generally level, with no significant deviation). This case is visualized in Fig. 2, for a coordinate $(u,v)$ in the attention feature space of one view, the corresponding epipolar line in the attention feature space of other views can be approximated as $Y=v$. This enables the extension of self-attention layers calculated on tokens within the same row across multiple views to learn 3D correspondences. As ablated in Tab. 4, the multi-view row attention can efficiently encourage view consistency with minor memory consumption.
Near-view Epipolar Aggregation. Though multi-view attention can effectively facilitate view consistency, we observe that the attention-only operation still struggles with accurate correspondences across views. To address this issue, we incorporate explicit feature aggregation among neighboring views to ensure multi-view consistency. Specifically, given the output features $\{{\mathbf{f}}_{\mathbf{v}}\}_{{\mathbf{v}}=1}^{N}$ from the multi-view row attention layers for each posed multi-view input, we propagate features by finding near-view correspondences with epipolar line constraints. Formally, for the feature map ${\mathbf{f}}_{{\mathbf{v}}}$ corresponding to the posed image ${\mathbf{x}}_{{\mathbf{v}}}$, we calculate its correspondence map $M_{{\mathbf{v}},{\mathbf{k}}}$ with a near-view ${\mathbf{x}}_{{\mathbf{k}}}$ as follows:
$M_{{\mathbf{v}},{\mathbf{k}}}[i]=\operatorname*{arg\,min}_{j,\ j{{}^{\top}}Fi=0}D({\mathbf{f}}_{\mathbf{v}}[i],{\mathbf{f}}_{{\mathbf{k}}}[j]),$ (2)
where $D$ denotes the cosine distance, and ${\mathbf{k}}\in\{{\mathbf{v}}-1,{\mathbf{v}}+1\}$ represents the two nearest neighbor views of the given pose. Here, $i$ and $j$ are indices of the spatial locations in the feature maps, $F$ is the fundamental matrix relating the two views ${\mathbf{v}}$ and ${\mathbf{k}}$, and the index $j$ lies on the epipolar line in the view ${\mathbf{k}}$, subject to the constraint $j^{\top}Fi=0$. We then obtain the aggregated feature map $\widetilde{{\mathbf{f}}}_{\mathbf{v}}$ of the view ${\mathbf{v}}$ by linearly combining features of correspondences from the two nearest views via:
$\widetilde{{\mathbf{f}}}_{\mathbf{v}}[i]$ $=w\cdot{\mathbf{f}}_{{\mathbf{v}}{-}1}[M_{{\mathbf{v}},{\mathbf{v}}-1}[i]]$ (3)
$\quad+(1-w)\cdot{\mathbf{f}}_{{\mathbf{v}}{+}1}[M_{{\mathbf{v}},{\mathbf{v}}+1}[i]],$
where $w$ represents the weight to combine the features of the two nearest views. The calculation of $w$ uses a hybrid fusion strategy, which ensures that the weight assignment accounts for both the physical camera distance and the token feature similarity (see the Appendix Sec. A.3). As the feature aggregation process is non-differentiable, we adopt the straight-through estimator $\text{sg}[\cdot]$ in VQVAE [63] to facilitate gradient back-propagation in the token space. Near-view epipolar aggregation explicitly propagates tokens from neighboring views, which greatly improves view consistency. However, due to substantial view changes, the corresponding tokens may not be available, leading to unexpected artifacts during token replacement. To address this, we fuse the feature ${\mathbf{f}}_{\mathbf{v}}$ of the current view with the feature $\widetilde{{\mathbf{f}}}_{\mathbf{v}}$ from near-view epipolar aggregation, with 0.5 averaging. This effectively combines multi-view row attention and near-view epipolar aggregation, thereby enhancing view consistency both implicitly and explicitly.
This approach is similar to token-space editing methods like TokenFlow [18] and DGE [11]. However, we propose a trainable version that considers both geometric and feature similarity for effective feature fusion.
3.3 Multi-view Data Augmentation
Our goal is to train a versatile and robust enhancement model that performs well on low-quality multi-view images from diverse data sources, such as those generated by image-to-3D models or rendered from coarse 3D representations. To achieve this, we carefully design a comprehensive data augmentation pipeline to expand the distribution of distortions in our base training data, bridging the domain gap between training and inference.
Texture Distortion. To emulate the low-quality textures and local inconsistencies found in synthesized multi-view images, we employ a texture degradation pipeline commonly used in 2D enhancement [69, 91]. This pipeline randomly applies downsampling, blurring, noise, and JPEG compression to degrade the image quality. To simulate inconsistencies between views, we adopt different stochastic degradation parameters for each view.
Texture Deformation and Camera Jitter. As in LGM [58], we introduce grid distortion to simulate texture inconsistencies in multi-view images and apply camera jitter augmentation to introduce variations in the conditional camera poses of multi-view inputs.
Color Shift. We also observe color variations in corresponding regions between multi-view images generated by image-to-3D models. By randomly applying color changes to some image patches, we encourage the model to produce results with consistent colors. In addition, renderings from a coarse 3DGS sometimes result in a grayish overlay or ghosting artifacts, akin to a translucent mask. To simulate this effect, we randomly apply a semi-transparent object mask to the image, allowing the model to learn to remove the overlay and improve 3D visual quality.
Noise-level Control. To control the enhancement strength, we apply noise augmentation by adding controllable noise to the input multi-view images. This noise augmentation process is similar to the diffusion process in diffusion models. This approach can further enhance the model’s robustness in handling unseen artifacts [91]. Generally, higher noise levels lead to stronger refinement and regeneration. This design allows users to adjust the noise levels, balancing the enhancement and generation, as shown in Fig. 1(b).
Figure 3: Qualitative comparisons of enhancing multi-view synthesis on the Objaverse synthetic dataset. As can be seen, only 3DEnhancer can correct flowed and missing textures with view consistency. (Zoom-in for best view)
3.4 Inference for 3D Enhancement
We present two ways to utilize our 3DEnhancer for 3D enhancement:
  • The proposed method can be directly applied to generation results from existing multi-view diffusion models [54, 39, 41, 25], and the enhanced output shall serve as the input to the multi-view 3D reconstruction models [58, 22, 71, 80]. Given the enhanced multi-view inputs with sharper textures and view-consistent geometry, our method can be directly used to improve the quality of existing multi-view to 3D reconstruction frameworks.
  • Our method can also be used for directly enhancing a coarse 3D model through iterative optimization. Specifically, given an initial coarse 3D reconstruction as $\mathcal{M}$ and a set of viewpoints $\{\pi_{\mathbf{v}}\}_{{\mathbf{v}}=1}^{N}$, we first render the corresponding views $\mathcal{X}=\{{\mathbf{x}}_{\mathbf{v}}\}_{{\mathbf{v}}=1}^{N}$ where ${\mathbf{x}}_{\mathbf{v}}=\text{Rend}(\mathcal{M},\pi_{\mathbf{v}})$ is obtained with the corresponding rendering techniques [43, 32]. Let $\mathcal{X}^{\prime}=\{{\mathbf{x}}_{\mathbf{v}}^{\prime}\}_{{\mathbf{v}}=1}^{N}$ be the enhanced multi-view images, we can then update the 3D model $\mathcal{M}$ by supervising it with $\mathcal{X}^{\prime}$ as
    $\mathcal{M}^{\prime}=\operatorname*{arg\,min}_{\mathcal{M}}\sum_{{\mathbf{v}}=1}^{N}\mathcal{L}({{\mathbf{x}}_{\mathbf{v}}^{\prime},\text{Rend}(\mathcal{M},\pi_{\mathbf{v}}))}.$ (4)
    Following previous methods that reconstruct 3D from synthesized 2D images [73, 17], we use a mixture of $\mathcal{L}_{\text{1}}$ and $\mathcal{L}_{\text{LPIPS}}$ [85] for robust optimization. In practice, unlike iterative dataset updates (IDU) [20], we found that inferring the enhanced views $\mathcal{X}^{\prime}$ once already yields high-quality results. More implementation details and results for this part are provided in the Appendix Sec. D.2.
4 Experiments
4.1 Datasets and Implementation
Figure 4: Qualitative comparisons of enhancing multi-view synthesis with RealBasicVSR[7] and Upscale-A-Video[91] on the in-the-wild dataset. Visually inspecting, 3DEnhancer yields sharp and consistent textures with intact semantics, such as the eyes of the girl in the mid row. (Zoom-in for best view)
Datasets. For training, we use the Objaverse dataset [14], specifically leveraging the G-buffer Objaverse [47], which provides diverse and high-quality renderings on Objaverse instances. We construct LQ-HQ view pairs following the augmentation pipeline outlined in Sec. 3.3 and then split the dataset into separate training and testing sets. Overall, approximately $400$K objects are used for training.
For evaluation, we use a test set containing $500$ objects from different categories within our synthesized Objaverse datasets. For each object, we render four orthogonal views from elevation angles ranging from $-5^{\circ}$ to $30^{\circ}$. Besides, we evaluate our model on the zero-shot in-the-wild dataset by selecting images from the GSO dataset [16], the output of generative image diffusion models [48], and web-sourced content. These images are then processed using several novel view synthesis methods [39, 54, 41, 36] to create our in-the-wild testing set, containing a total of 400 instances.
Implementation Details. We employ PixArt-$\Sigma$ [10], an efficient DiT model, as our backbone. Our model is trained on images with a resolution of 512 x 512. The AdamW optimizer [33] is used with a fixed learning rate of 2e-5. Our training is conducted over 10 days using 8 Nvidia A100-80G GPUs, with a batch size 256. For inference, we employ a DDIM sampler [56] with 20 steps and set the Classifier-Free Guidance (CFG) [23] scale to 4.5. During training, we randomly sample input views with azimuth angles ranging from $0^{\circ}$ to $360^{\circ}$ and elevation angles between $-5^{\circ}$ and $30^{\circ}$.
Baselines. To assess the effectiveness of our approach, we adopt two image enhancement models, RealESRGAN [69] and StableSR [64], along with two video enhancement models, RealBasicVSR [7] and Upscale-a-Video [91] as our baselines. For a fair comparison, we further fine-tune all these methods on the Objaverse dataset to minimize potential data domain discrepancies. During inference, since Real-ESRGAN, RealBasicVSR, and Upscale-a-Video by default produce images upscaled by a factor of ×4, we resize their outputs to the a uniform resolution of 512 × 512 for comparison.
Metrics. We evaluate the effectiveness of our methods on two tasks: multi-view synthesis enhancement and 3D reconstruction improvement. We employ standard metrics including PSNR, SSIM, and LPIPS [85] on our synthetic dataset, along with non-reference metrics including FID [51], Inception Score [49], and MUSIQ [31] on the in-the-wild dataset. For FID computation, we use the rendered images from Objaverse to represent the real distribution. Besides the quantitative metrics, we present comprehensive qualitative results in the paper.
Table 1: Quantitative comparisons of enhancing multi-view synthesis on the Objaverse synthetic dataset, the best and second-best results are marked in red and blue, respectively.
[…]
Input 26.15 0.9056 0.1257
Real-ESRGAN [69] 26.02 0.9185 0.0877
StableSR [64] 25.12 0.8914 0.1130
RealBasicVSR [7] 26.21 0.9212 0.0888
Upscale-A-Video [91] 25.57 0.8937 0.1153
3DEnhancer (Ours) 27.53 0.9265 0.0626
4.2 Comparisons
Table 2: Quantitative comparisons of enhancing multi-view synthesis on the in-the-wild dataset.
[…]
Generated MV 52.77 112.12 7.68 $\pm$ 0.86
Real-ESRGAN [69] 72.47 114.25 7.31 $\pm$ 0.89
StableSR [64] 70.43 111.53 7.59 $\pm$ 0.97
RealBasicVSR [7] 74.07 128.30 7.09 $\pm$ 0.87
Upscale-A-Video [91] 71.73 114.81 7.75 $\pm$ 0.96
3DEnhancer (Ours) 73.32 108.40 7.93 $\pm$ 1.11
Enhancing Multi-view Synthesis. The output images from multi-view synthesis models often lack texture details or exhibit inconsistencies across views, as shown in Fig. 1. To demonstrate that 3DEnhancer can correct flawed textures and recover missing textures, we provide quantitative results on both the Objaverse synthetic dataset and the in-the-wild dataset in Tab. 1 and Tab. 2, respectively. Qualitative comparisons on both test sets are presented in Fig. 3 and Fig. 4. As can be seen, our method outperforms others across most metrics. While RealBasicVSR achieves a higher MUSIQ score on the in-the-wild dataset, it fails to generate visually plausible images, as shown in Fig. 4. The image enhancement models RealESRGAN and StableSR can recover textures to some extent in individual views, but they fail to maintain consistency across multiple views. Video enhancement models, such as RealBasicVSR and Upscale-A-Video, also fail to correct texture distortions effectively. For example, both models fail to generate smooth facial textures in the first example shown in Fig. 4. In contrast, our method generates more natural and consistent details across views.
Table 3: Quantitative comparisons of enhancing 3d reconstruction on the in-the-wild dataset.
[…]
Input 41.04 77.54 8.98 $\pm$ 0.65
Real-ESRGAN [69] 65.25 74.29 8.29 $\pm$ 0.35
StableSR [64] 65.71 72.98 9.51 $\pm$ 0.78
RealBasicVSR [7] 65.70 74.58 7.77 $\pm$ 0.48
Upscale-A-Video [91] 64.05 74.12 9.77 $\pm$ 0.79
3DEnhancer (Ours) 66.04 71.78 9.96 $\pm$ 0.96
Figure 5: Qualitative comparisons of enhancing 3D reconstruction given generated multi-view images on the in-the-wild dataset. Due to the low quality and view-inconsistency of outputs from multi-view models, the reconstructed 3D models often suffer from issues such as blurry textures, like the skull’s teeth in the first row, or artifacts such as ghosting hands in the second Mario example. Existing enhancement methods fail to correct texture artifacts. In contrast, our method produces both geometrically accurate and visually appealing results. (Zoom-in for best view)
Figure 6: Low-resolution 3D Gaussians optimization with 3DEnhancer. (Zoom-in for best view)
Enhancing 3D Reconstruction. In this section, we present 3D reconstruction comparisons based on rendering views from 3D Gaussians generated by LGM [58]. Quantitative comparisons are shown in Tab. 3. The proposed method outperforms previous approaches in terms of 3D reconstruction. For qualitative evaluation, we visualize the results of two video enhancement models, RealBasicVSR and Upscale-A-Video. As shown in Fig. 5, these baselines suffer from a lack of multi-view consistency, leading to misalignment, such as the misalignment of the teeth in the first skull example and the ghosting in the example of Mario’s hand. In contrast, our model maintains consistency and produces high-quality texture details. We further demonstrate our approach can optimize a coarse differentiable representation. As shown in Fig. 6, our method is capable of refining low-resolution Gaussians [52]. More details and results of refining coarse Gaussians are provided in the Appendix Sec. D.2.
Table 4: Ablation study of cross-view modules.
[…]
(a) 25.11 0.9067 0.081
(b) 25.95 0.9147 0.072
(c) 26.92 0.9226 0.0642
(d) 27.53 0.9265 0.0626
4.3 Ablation Study
Effectiveness of Cross-View Modules. To evaluate the effectiveness of our proposed cross-view modules, we ablate two modules: multi-view row attention and near-view epipolar aggregation. As shown in Tab. 4, removing either module results in worse textures between views. The visual comparison in Fig. 7 also validates this observation. Without the multi-view row attention module, the model fails to produce smooth textures, as shown in Fig. 7 (b). Without the epipolar aggregation module, reduced texture consistency is observed, as depicted in Fig. 7 (c).
Besides, the epipolar constraint is essential for preventing the model from learning textures from incorrect regions in other views and contributes to the overall consistency. As demonstrated in Fig. 8, without the epipolar constraint, the texture of the top part of the flail is incorrectly aggregated from the grip in the other view, thus resulting in inconsistency across views.
Figure 7: Effectiveness of cross-view modules.
Figure 8: Comparisons of enhancing multiview images with and without epipolar aggregation. The red line denotes the epipolar line corresponding to the circled area, while the dotted arrow indicates the target location of tokens propagated from one view to another. (Zoom-in for best view)
Effectiveness of Noise Level. As shown in Fig 1, our model can generate diverse textures by adjusting noise levels. Low noise levels generally result in outputs with blurred details, while high noise levels produce sharper, more detailed textures. However, high noise levels may also reduce the fidelity of the input images.
5 Conclusion
In conclusion, this work presents a novel 3D enhancement framework that leverages view-consistent latent diffusion model to improve the quality of given coarse multi-view images. Our approach introduces a versatile pipeline combining data augmentation, multi-view attention and epipolar aggregation modules that effectively enforces view consistency and refines textures across multi-view inputs. Extensive experiments and ablation studies demonstrate the superior performance of our method in achieving high-quality, consistent 3D content, significantly outperforming existing alternatives. This framework establishes a flexible and powerful solution for generic 3D enhancement, with broad applications in 3D content generation and editing.
Appendix
In this appendix, we provide additional discussions and results to supplement the main paper. In Sec. A, we present more architecture and design details of our 3DEnhancer. In Sec. B, we provide detailed information about our training dataset, including the augmentation pipeline and illustrative examples. Sec. C highlights some interesting findings related to inference. Comprehensive results are presented in Sec. D to further demonstrate our performance. Notably, we also include a demo video (Sec. D.4) to showcase rendering results for 3D reconstruction enhancement.
A Architecture and Design
A.1 Pose-aware Encoder
Our pose-aware encoder is adapted from the convolutional encoder of LDM [48]. As shown in Fig. 2, the output of the pose-aware encoder serves as the conditioning features for the trainable copies in our ControlNet [83]. The details of its hyperparameters are summarized in Tab. 5. This encoder employs 64 channels and a single residual block to enhance efficiency. Additionally, we incorporate cross-view self-attention [54] into the middle layer of the encoder to improve inter-view consistency. To ensure compatibility with the number of latent channels in the DiT blocks, the output $z$-channels number is set to 1152. The final convolutional layer in the encoder uses a stride of 2 to match the dimensions of the DiT block latents. All other hyperparameters are kept at default values.
A.2 View-Consistent DiT Block
The view-consistent DiT block is based on the PixArt-$\Sigma$ [10] architecture. Consistent with PixArt-$\Sigma$, we use the T5 large language model as the text encoder for conditional text feature extraction, and the frozen VAE from SDXL [45] to capture the latent features of images. PixArt-$\Sigma$ consists of 28 Transformer blocks. For the ControlNet [83] implementation, we utilize trainable copies of the first 13 base blocks, augmenting each copied block with zero linear layers before and after it. The output of the $i$-th trainable copied block is added to the corresponding frozen base $i$-th block. The multi-view row attention with near-view epipolar aggregation is an additional attention layer that is inserted into both the DiT blocks and the copied ControlNet blocks. This layer is positioned after the self-attention layer, as illustrated in Fig. 2. During training, we train the entire ControlNet blocks and every inserted multi-view row attention layer in the DiT blocks. Detailed hyperparameters for the DiT block and the inserted row attention layers are provided in Tab. 5.
A.3 Weight for Two Nearest Views Aggregation
In Eq. 3, we compute the fusion weight $w$ based on both the physical camera distance and the similarity of token features. First, we consider the geometric distance weight $w_{d}$, which reflects the proximity of the camera:
$w_{d}=\frac{d_{{\mathbf{v}},{\mathbf{v}}+1}}{d_{{\mathbf{v}},{\mathbf{v}}-1}+d_{{\mathbf{v}},{\mathbf{v}}+1}},$ (5)
where $d_{{\mathbf{v}},{\mathbf{k}}}$ represents the geometric distance between the camera of view $v$ and the camera of view ${\mathbf{k}}\in\{{\mathbf{v}}-1,{\mathbf{v}}+1\}$. To ensure the nearest-view weight calculation also incorporates token feature similarity, we augment the weight token-wise with token similarity:
$w=\frac{S^{i}_{{\mathbf{v}},{\mathbf{v}}-1}\cdot w_{d}}{S^{i}_{{\mathbf{v}},{\mathbf{v}}-1}\cdot w_{d}+(1-w_{d})\cdot S^{i}_{{\mathbf{v}},{\mathbf{v}}+1}},$ (6)
where $S^{i}_{{\mathbf{v}},{\mathbf{k}}}$ denotes the cosine similarity of the corresponding tokens, i.e., ${\mathbf{f}}_{\mathbf{v}}[i]$ and ${\mathbf{f}}_{{\mathbf{k}}}[M_{{\mathbf{v}},{\mathbf{k}}}[i]]$.
Table 5: Hyperparameters for the pose-aware encoder, view-consistent DiT block, and the inserted multi-view row attention layers in our 3DEnhancer. The table follows the hyperparameter table style from [48, 44]. We train our model on images with a resolution of $512\times 512$ using 4 views.
Hyperparameter DiT
Layers 28
Training views shape $4\times 512\times 512\times 3$
$f$ 8
Patch size 2
Embedding dimension 1024
Hidden size 1152
$z$-shape $4\times 1024\times 1152$
Head number 16
CA sequence length 300
Hyperparameter Pose-aware Encoder
$f$ 8
Channels 64
Channel multiplier 1, 2, 4, 4
$z$-channels 1152
Hyperparameter Row Attention
Head number 16
Positional encoding sine-cosine
Epipolar aggregation True
B Dataset
B.1 Dataset
The G-buffer Objaverse dataset [47] contains a broad variety of 3D objects categorized into 10 types: Human-Shaped, Animals, Daily Objects, Furniture, Buildings and Outdoor Objects, Transportation, Plants, Food, and Electronics. To ensure high standards, we exclude any objects labeled as “Poor-quality.” We observe that the original captions in G-buffer Objaverse are simple and lack detailed information. Therefore, we adopt captions from 3D-Topia [26], which provide more informative and accurate descriptions for a subset of objects in Objaverse. We update the caption of each object accordingly if it exists in 3D-Topia, resulting in the refinement of approximately 45% of the captions. Additionally, to facilitate CFG [23], we omit the text condition at a rate of 0.2. Such settings enhance the robustness of our method to text conditions with varying levels of detail. For the in-the-wild dataset, we remove backgrounds and center objects as previous works [65, 41, 36]. We uniformly apply a white background to the input views.
Figure 9: Visualization of several examples from our augmentation pipeline. Thanks to the comprehensive augmentation strategy, our method is able to bridge the domain gap between training and inference.
B.2 Data Augmentation
The visualization of the data augmentation pipeline is shown in Fig. 9. During training, we dynamically generate synthetic training pairs on the fly, and the argumentation is implemented in PyTorch with CUDA acceleration to ensure efficiency. The pipeline incorporates several stochastic augmentation steps, producing diverse training pairs with varying levels of degradation. During augmentation, the input views of the same object are either augmented with the same level of degradation (e.g., the same blur kernel) or with different stochastic augmentations. This strategy encourages the model’s ability to learn information across views, particularly from those with fewer degradations. We ensure that the augmentation is confined to the object’s masked area with a slight mask dilation. This allows the white background unaffected, which aligns with real-world scenarios of low-quality multi-view images. We also set a probability where no augmentation is applied to the input images, i.e., the low-quality images are identical to the ground truth. In such cases, the model is encouraged to preserve fidelity when the input images are already of high-quality. Details of several augmentation parameters are summarized in Tab. 6. Further implementation details will be provided in our code release.
Table 6: Several augmentation parameters that are used in our augmentation pipeline.
Argumentation type Parameters
First-order blur prob 0.8
Second-order blur prob 0.3
Blur kernel size range {7, 9, …, 21}
Blur standard deviation range [0.2, 3]
Gaussian noises prob 0.5
Resize range [0.3, 1.5]
JPEG compression quality factor [80, 100]
Argumentation type Parameters
Final sinc filter prob 0.8
Camera jitter prob 0.2
Camera jitter strength range [0.05, 0.1]
Color shift prob 0.3
Grid distortion prob 0.3
Grid distortion strength range [0.2, 0.5]
No argumentation prob 0.1
C More Details on Inference
C.1 Multi-View Editing
Benefiting from our comprehensive augmentation pipeline and the robust view-consistent DiT Block, we observe an interesting fact: our method is capable of generating detailed and consistent textures even from extremely coarse or corrupted multi-view inputs. This enables our approach to modify multi-view images in two distinct ways: 1. Applying a black mask to the region designated for editing and modifying the text prompt to generate the target multi-view images. 2. Adjusting the inference noise level, where higher noise levels produce more diverse outputs. Using the edited multi-view images, we can subsequently modify the reconstructed 3D representations. An example of editing 3D Gaussians generated by LGM [58] through modifying its multi-view input is shown in Fig. 10.
Figure 10: Rendered views of edited 3D Gaussians using our multi-view editing approach. By adding a large noise or a black mask, and leveraging text prompts as guidance, we consistently modify the texture of the bags.
C.2 Color Correction
Previous studies [64, 91] have highlighted that diffusion models often exhibit color shift artifacts, where the global color scheme deviates from the input images. This is different from our color shift augmentation, which introduces localized color changes to specific image regions. However, this augmentation also aims to encourage the model to maintain consistent color reproduction. We observe that integrating a training-free wavelet color correction module [64] can help resolve the global color scheme shift. As reported in Tab. 8, applying wavelet color correction leads to improved fidelity metrics (higher PSNR, SSIM, and lower LPIPS [85]) for the baseline, but it has minimal impact on our results, showing our robustness against global color scheme shifts. However, at extremely high noise levels, such as $\delta=200$, minor global color shifts may still occur in our method because the noise may impact the original color information. In such cases, wavelet color correction could be beneficial, as illustrated in Fig. 11.
Figure 11: Minor global color scheme shift at high noise levels. When the noise level $\delta$ is small, such as $\delta=0$, our method maintains excellent color fidelity. However, at a higher noise level, such as $\delta=200$ in the example, the output figure’s face appears slightly darker than that of the input. In this case, the wavelet color correction [64] could help mitigate this issue.
D More Results
D.1 User Study
To enable a thorough comparison, we conduct a user study to evaluate the enhancement results of multi-view images and 3D reconstructions. For the multi-view image enhancement, each participant is shown 10 sets of randomly selected objects’ multi-view images, enhanced by our 3DEnhancer, RealESRGAN [69], StableSR [64], RealBasicVSR [7], and Upscale-a-Video [91]. For the 3D reconstruction enhancement, participants are presented with another 10 360-degree rotating render videos of the 3D Gaussians enhanced by our method, RealBasicVSR [7], and Upscale-a-Video [91]. Their task is to choose the visually superior enhanced results. A total of 20 participants take part in the study. As illustrated in Fig. 12, The results indicate a strong preference for our method over the compared approach. On average, 74% of users preferred our method for enhancing multi-view images, while 78% favored it for enhancing 3D reconstruction. These findings strongly demonstrate the quality and robustness of our approach.
Figure 12: User study results. Human voters consistently prefer our method over other approaches.
D.2 Results of Optimizing 3D Gaussians
3D representations can be rendered from multiple views, this nature allows our method to iteratively optimize a coarse 3D representations. To demonstrate this capability, we adopt Gaussian Splatting [32] as our example due to its high rendering fidelity and efficiency. Specifically, we implement a pipeline to refine coarse 3D Gaussians checkpoints by leveraging our enhanced outputs as pseudo ground truth. We randomly select 20 objects from the Objaverse test dataset for evaluation. Following [52], we fit low-resolution 3D Gaussians using images obtained by bilinearly downsampling the original dataset images by a factor of 8, resulting in a resolution of 64 × 64 pixels. We use three distinct trajectories for fitting low-resolution Gaussians, refining Gaussians, and evaluation. As proposed in [17], our refinement process also minimizes a combined loss function, including a photometric reconstruction loss and a perceptual loss [85]. The perceptual loss emphasizes high-level semantic similarity between rendered and enhanced images while ignoring inconsistencies in low-level, high-frequency details. To improve regularization during refining, we sample 100 views along a single smooth orbital path, as increasing the number of views has been shown to enhance the refining process [17]. The optimization is conducted over 2000 refinement steps for all methods and takes approximately 130s to refine a single object on one NVIDIA A100 GPU. For comparison, we evaluate our method against two video enhancement models, RealBasicVSR  [7] and Upscale-A-Video [91]. Quantitative and qualitative results are presented in Tab. 7 and Fig. 13, respectively. Our results demonstrate detailed and sharp outputs, while other methods exhibit ghosting artifacts and blurry textures. The results highlight the superior performance of our approach in refining coarse 3D representations.
Table 7: Quantitative comparison of optimizing low-resolution Gaussians. The best results are highlighted in bold.
Metrics Low-Resolution Gaussians RealBasicVSR [7] Upscale-A-Video[91] 3DEnhancer
PSNR $\uparrow$ 26.35 27.39 26.20 27.54
SSIM $\uparrow$ 0.9120 0.9216 0.9184 0.9337
LPIPS $\downarrow$ 0.1135 0.0803 0.0928 0.0756
Figure 13: Qualitative comparisons of optimizing low-resolution Gaussians. During optimization, both RealBasicVSR [7] and Upscale-A-Video  [91] produce ghosting and blurry textures due to inconsistent outputs. Our 3DEnhancer achieves sharp and clear results.
D.3 More Comparisons
In this section, we introduce another baseline from the multi-view image upscale module in Unique3D [72]. This baseline finetunes ControlNet-Tile [83] to enhance RGB views. While the module can sharpen some textures, it struggles to recover inconsistent or corrupted areas in multi-view images. Our method outperforms Unique3D’s MV Upscale both quantitatively and qualitatively. The quantitative comparison between Unique3D’s MV Upscale and our method is presented in Tab. 8. Additionally, we provide more visual comparisons of our method with all other baselines, including RealESRGAN [69], StableSR [64], Unique3D’s MV Upscale [72], RealBasicVSR [7], and Upscale-a-Video [91]. Fig. 14 and Fig. 15 showcase the visual comparisons of multi-view enhancement on synthetic and in-the-wild datasets, respectively.
Table 8: Quantitative comparison of enhancing multi-view synthesis on the Objaverse synthetic dataset with Unique3D’s MV Upscale module. Our method demonstrates clear advantages in restoration fidelity, as measured by PSNR, SSIM, and LPIPS. While applying color correction improves the output of Unique3D’s MV Upscale module, it has minimal impact on our results when noise level is set to 0, highlighting our method’s robustness against global color scheme shift issues.
Metrics Unique3D’s Upscale Unique3D’s Upscale (+ color fix) 3DEnhancer 3DEnhancer (+ color fix)
PSNR $\uparrow$ 25.75 26.18 27.53 27.50
SSIM $\uparrow$ 0.8989 0.9055 0.9265 0.9258
LPIPS $\downarrow$ 0.1300 0.1257 0.0626 0.0631
Figure 14: Qualitative comparisons on the Objaverse synthetic dataset. Our 3DEnhancer demonstrates promising improvements, with increased detail and enhanced realism. (Zoom in for best view.)
Figure 15: Qualitative comparisons on the in-the-wild dataset. Our 3DEnhancer yields significant improvements, providing enhanced detail and consistent output. (Zoom in for best view.)
D.4 Video Demo
We also provide a demo video (3DEnhancer-demo.mp4) in our project page, showcasing visual results of 3D reconstruction enhancement.